OpenStudy (anonymous):

How can you put 2(x+1)+0 into the quadratic equation?

OpenStudy (abb0t):

distribute the 2 across the parenthesis would be your first step.

OpenStudy (anonymous):

I have that, but it doesn't look right, lol.

OpenStudy (abb0t):

\[ax^2 +bx +c\] where quadratic formulas: \[\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (abb0t):

If you don't have an \[ax^2\] then that must mean that a = 0, right? Since anything multiplied by zero = zero.

OpenStudy (abb0t):

But the problem doesn't look right to me either.

OpenStudy (anonymous):

So maybe there is no ax^2?

OpenStudy (mertsj):

That is not a quadratic equation. Why would you use the quadratic formula?

OpenStudy (abb0t):

@Mertsj is right. You wouldn't use it, but theoretically, right? wouldn't you assume zero for a?

OpenStudy (anonymous):

I'n trying to find the graph with a discriminant of 0 and what i gave is the lines equation.

OpenStudy (mertsj):

Yes. If you wanated to put it into quadratic form, a would have to be 0

OpenStudy (anonymous):

Okay, thanks. It really confused me. But thanks to both of you!

OpenStudy (mertsj):

0x^2+2x + 2 = 0